An unknown gas occupies a volume of 4.75 L at 1500 K and 5.00 atm. If its mass is 5.45 g, what is the molar mass of the gas?


1
Expert's answer
2021-05-06T07:39:03-0400

Volume of the gas = 4.75L

Temperature= 1500K

Pressure= 5.00atm

no of moles = ?

Rate = 0.082


To calculate the molar mass, the no of moles must be calculated first by using the Gas law equation

PV = nRT

n = PV/RT

n= 5 x 4.75/0.082 x 1500

= 23.75/123

= 0.193mol

The number of moles = 0.193mol

Mass given = 5.45g

No of moles = 0.193mol

Molar Mass= Mass given/ no of moles

Molar mass = 5.45/0.193

Molar Mass= 28.24g/mol

Hence, the Molar Mass of the unknown gas is 28.24g/mol